OpenStudy (anonymous):

((4/x-5)+(2/x+2))/((2x/(x^2-3x-10))+3) simplify and show steps

OpenStudy (anonymous):

\[\frac{\frac{4}{x-5}+\frac{2}{x+2}}{\frac{2x}{x^2-3x-10}+3}???\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the question but you cant do that in the ask bar

OpenStudy (anonymous):

\[\frac{\frac{4(x+2)+2(x-5)}{(x-5)(x+2)}}{\frac{2x+3(x^2-3x-10)}{x^2-3x-10}}\] \[\frac{4x+8+2x-10}{x^2+2x-5x-10}*\frac{x^2-3x-10}{2x+3x^2-9x-30}\] \[\frac{6x-2}{\cancel{x^2-3x-10}}*\frac{\cancel{x^2-3x-10}}{3x^2-7x-30}\] \[\frac{6x-2}{3x^2-7x-30}\]